OpenStudy (anonymous):

F(X)=3X-6+g(X)=X-2 FIND F/G AND IT'S DOMAIN.

OpenStudy (anonymous):

g(x)=4-2x f(x)=x-2 f/g(x)=(x-2)/(4-2x) domain =R-{2}

OpenStudy (anonymous):